OpenStudy (highschoolmom2010):

find geometric mean of 1/2,2

OpenStudy (rsadhvika):

multiply them, take square root

OpenStudy (cwrw238):

geometric mean of x and y is sqrt(x* y)

OpenStudy (highschoolmom2010):

.5*2=1 \[\sqrt{1}=1\]

OpenStudy (cwrw238):

right

OpenStudy (highschoolmom2010):

so its 1 that was the dumbest problem ive seen yet

OpenStudy (rsadhvika):

\(\large \frac{1}{2}\times 2\) \(\large \frac{1}{\cancel{2}}\times \cancel{2}\) \(\large \sqrt{1}\) \(\large 1\)

OpenStudy (rsadhvika):

no problem is dumbest in math

OpenStudy (cwrw238):

its not really dumb

OpenStudy (rsadhvika):

cuz there will be another problem thats more dumber than this one lol... like we dont have a largest number

OpenStudy (highschoolmom2010):

lol i just dont see why they had to make me find sqrt of 1

OpenStudy (rsadhvika):

you got lucky today.... :P

OpenStudy (rsadhvika):

if u look at it more carefully, geometric mean of ANY number and its reciprocal is always 1

OpenStudy (highschoolmom2010):

yepp

OpenStudy (rsadhvika):

isnt that complex enough to ponder over and hurt our brains

OpenStudy (highschoolmom2010):

nope not yet

OpenStudy (rsadhvika):

why geometric mean of any number and its reciprocal equals 1 ?